OpenStudy (anonymous):

HELP PLEASE!!! What is the solution set of the equation 2 cos^2 data- cos data =0 over the interval 0 degrees < or equal to data < 360 degrees?

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

http://en.wikipedia.org/wiki/Data_(Star_Trek)

OpenStudy (misty1212):

start with \[2\cos(\theta)(\cos(\theta)-1)=0\]

OpenStudy (misty1212):

then solve \[\cos(\theta)=0\] and also \[\cos(\theta)-1=0\]

OpenStudy (anonymous):

the first one is pie over 2 the second also pie over 2 I think